# SEEK: Self-adaptive Explainable Kernel For Nonstationary Gaussian Processes

### v1

@@ Abstract
Gaussian processes (GPs) are powerful probabilistic models that define flexible priors over functions, offering strong interpretability and uncertainty quantification. However, GP models often rely on simple, stationary kernels which can lead to suboptimal predictions and miscalibrated uncertainty estimates, especially in nonstationary real-world applications. In this paper, we introduce SEEK, a novel class of learnable kernels to model complex, nonstationary functions via GPs. Inspired by artificial neurons, SEEK is derived from first principles to ensure symmetry and positive semi-definiteness, key properties of valid kernels. The proposed method achieves flexible and adaptive nonstationarity by learning a mapping from a set of base kernels. Compared to existing techniques, our approach is more interpretable and much less prone to overfitting. We conduct comprehensive sensitivity analyses and comparative studies to demonstrate that our approach is not robust to only many of its design choices, but also outperforms existing stationary/nonstationary kernels in both mean prediction accuracy and uncertainty quantification.
@@ 2.1 Gaussian Processes (GPs)
$\bar{m}\mathopen{}\left(\boldsymbol{x}^{*}\right)\mathclose{}=m\mathopen{}\left(\boldsymbol{x}\right)\mathclose{}+c\mathopen{}\left(\boldsymbol{x}^{*},\boldsymbol{X}\right)\mathclose{}\boldsymbol{C_{\lambda}}^{-1}\mathopen{}\left(\boldsymbol{y}-\boldsymbol{m}\right)\mathclose{},$ (2a)
@@ 2.4 SEEK: Our Proposed Kernel
Let $\boldsymbol{x},\boldsymbol{x}^{\prime}\in{}^{P}$ be two input points, and consider vector-valued functions $\boldsymbol{w}_{m}(\boldsymbol{x};\boldsymbol{\beta}_{\boldsymbol{w}_{m}}):{}^{P}\rightarrow{}^{W_{m}}$ and $\boldsymbol{b}(\boldsymbol{x};\boldsymbol{\beta}_{b}):{}^{P}\rightarrow{}^{B}$, parameterized via $\boldsymbol{\beta}_{\boldsymbol{w}_{m}}$ and $\boldsymbol{\beta}_{b}$, respectively. Given a set of $M$ base kernels $\mathopen{}\left\{c_{m}\mathopen{}\left(\boldsymbol{x},\boldsymbol{x}^{\prime};\boldsymbol{\beta}_{c_{m}}\right)\mathclose{}\right\}\mathclose{}_{m=1}^{M}$ and an appropriate activation function $\phi:\real\rightarrow\real$, we define SEEK as:
@@
$z(\boldsymbol{x},\boldsymbol{x}^{\prime})=\sum_{m=1}^{M}\boldsymbol{w}_{m}(\boldsymbol{x})\boldsymbol{w}_{m}^{\top}(\boldsymbol{x}^{\prime})c_{m}(\boldsymbol{x},\boldsymbol{x}^{\prime})+\boldsymbol{b}(\boldsymbol{x})\boldsymbol{b}^{\top}(\boldsymbol{x}^{\prime}).$ (6)
@@ 2.5 Explainability of SEEK: Illustrative Example
From Figure 2(a), we observe that our GP (1) produces highly accurate predictions, achieving an RMSE of $0.013$ which is notably smaller than $0.038$ reported in [noack2024unifying], and (2) provides well-calibrated prediction intervals. Figures 2(b) and 2(c) illustrate the learned weighted Gaussian and periodic kernels, respectively. The weighted Gaussian kernel exhibits a similar trend at both reference points where it has a maximum that is smoothly decreased. Unlike the Gaussian kernel, the weighted periodic kernel varies adaptively. Specifically, for $x^{\prime}=0.3$, the kernel exhibits little to no variation across the entire domain, aligning with the smooth nature of the underlying function. In contrast, at $x^{\prime}=0.9$, the kernel adjusts to local variations, i.e., it exhibits an oscillating pattern only in regions where the underlying function does as well. Finally, Figure 2(d) presents the resulting covariance function obtained by Eq. 5. At a large scale, the covariance functions for both reference points appear similar. However, a closer look (see insets) reveals clear distinct behaviors: $c(x,x^{\prime}=0.9)$ exhibits a more oscillatory structure, driven by the learned weighted periodic kernel, whereas $c(x,x^{\prime}=0.3)$ presents a smoother behavior.
@@ 3 Results and Discussions
All simulations in this section are repeated $16$ times to ensure the results are representative. Also, all the GPs are trained via the L-BFGS optimizer with a learning rate of $0.01$. For further implementation and optimization details, please see Section A in the Appendix.
@@ 3.1 Benchmark problems and metrics
We consider four benchmark problems with varying levels of nonlinearity. Some of these functions are adopted from the literature [chen2024adaptive, noack2024unifying, picheny2013benchmark].
@@ 3.2 Sensitivity Studies
We evaluate the performance of the models in terms of NRMSE and NNOIS. The results of these studies are summarized in Figure 3, where each row corresponds to a different study. The analyses are assessed on the Analytic II function and the Volcano dataset.
@@
This observation poses an interesting dilemma: while nonstationary kernels are more beneficial in low-to-mid data regimes, their increased flexibility also makes them more prone to overfitting compared to stationary kernels. This highlights the need for the design of nonstationary kernels with safeguards built into their structure that makes them robust against overparameterization.
@@
Figure 4: Our proposed kernel SEEK with (a) 1 and (b) 6 Gaussian kernels as base kernels on the Volcano dataset: Increasing the number of base kernels reduces error and improves confidence, as reflected in the colorbars. Black dots represent the training data points.
@@ 3.3 Comparative Studies
In our studies, we noticed that nonstationary kernels generally require more data to outperform the Gaussian kernel in high-dimensional settings where data becomes sparse and learning complex, input-dependent variations is more challenging. While nonstationary kernels introduce additional flexibility to capture such local variations, they also have a higher risk of overfitting when data is limited. In contrast, the Gaussian kernel imposes a globally smooth prior, acting as an effective regularizer. This observation aligns with prior works [remes2018neural, risser2020bayesian, noack2023exact], which combine nonstationary kernels with sparse approximation techniques to allow the GP to leverage large datasets in high dimensions.

### v2

@@ Abstract
Gaussian processes (GPs) are powerful probabilistic models that define flexible priors over functions, offering strong interpretability and uncertainty quantification. However, GP models often rely on simple, stationary kernels which can lead to suboptimal predictions and miscalibrated uncertainty estimates, especially in nonstationary real-world applications. In this paper, we introduce SEEK, a novel class of learnable kernels to model complex, nonstationary functions via GPs. Inspired by artificial neurons, SEEK is derived from first principles to ensure symmetry and positive semi-definiteness, key properties of valid kernels. The proposed method achieves flexible and adaptive nonstationarity by learning a mapping from a set of base kernels. Compared to existing techniques, our approach is more interpretable and much less prone to overfitting. We conduct comprehensive sensitivity analyses and comparative studies to demonstrate that our approach is not only robust to many of its design choices, but also outperforms existing stationary/nonstationary kernels in both mean prediction accuracy and uncertainty quantification.
@@ 2.1 Gaussian Processes (GPs)
$\bar{m}\mathopen{}\left(\boldsymbol{x}^{*}\right)\mathclose{}=m\mathopen{}\left(\boldsymbol{x}^{*}\right)\mathclose{}+c\mathopen{}\left(\boldsymbol{x}^{*},\boldsymbol{X}\right)\mathclose{}\boldsymbol{C_{\lambda}}^{-1}\mathopen{}\left(\boldsymbol{y}-\boldsymbol{m}\right)\mathclose{},$ (2a)
@@ 2.4 SEEK: Our Proposed Kernel
Let $\boldsymbol{x},\boldsymbol{x}^{\prime}\in{}^{P}$ be two input points, and consider vector-valued functions $\boldsymbol{w}_{m}(\boldsymbol{x};\boldsymbol{\beta}_{\boldsymbol{w}_{m}}):{}^{P}\rightarrow{}^{W_{m}}$ and $\boldsymbol{b}(\boldsymbol{x};\boldsymbol{\beta}_{b}):{}^{P}\rightarrow{}^{B}$, parameterized via $\boldsymbol{\beta}_{\boldsymbol{w}_{m}}$ and $\boldsymbol{\beta}_{b}$, respectively. [^2] Unless otherwise specified, all vectors are treated as column vectors. Given a set of $M$ base kernels $\mathopen{}\left\{c_{m}\mathopen{}\left(\boldsymbol{x},\boldsymbol{x}^{\prime};\boldsymbol{\beta}_{c_{m}}\right)\mathclose{}\right\}\mathclose{}_{m=1}^{M}$ and an appropriate activation function $\phi:\real\rightarrow\real$, we define SEEK as:
@@
$z(\boldsymbol{x},\boldsymbol{x}^{\prime})=\sum_{m=1}^{M}\boldsymbol{w}_{m}^{\top}(\boldsymbol{x})\boldsymbol{w}_{m}(\boldsymbol{x}^{\prime})c_{m}(\boldsymbol{x},\boldsymbol{x}^{\prime})+\boldsymbol{b}^{\top}(\boldsymbol{x})\boldsymbol{b}(\boldsymbol{x}^{\prime}).$ (6)
@@ 2.5 Explainability of SEEK: Illustrative Example
From Figure 2(a), we observe that our GP (1) produces highly accurate predictions, achieving an RMSE of $0.013$ which is notably smaller than $0.038$ reported in [noack2024unifying], and (2) provides well-calibrated prediction intervals, obtaining an NNOIS of $0.189$ (see Eq. 10b). Figures 2(b) and 2(c) illustrate the learned weighted Gaussian and periodic kernels, respectively. The weighted Gaussian kernel exhibits a similar trend at both reference points where it has a maximum that is smoothly decreased. Unlike the Gaussian kernel, the weighted periodic kernel varies adaptively. Specifically, for $x^{\prime}=0.3$, the kernel exhibits little to no variation across the entire domain, aligning with the smooth nature of the underlying function. In contrast, at $x^{\prime}=0.9$, the kernel adjusts to local variations, i.e., it exhibits an oscillating pattern only in regions where the underlying function does as well. Finally, Figure 2(d) presents the resulting covariance function obtained by Eq. 5. At a large scale, the covariance functions for both reference points appear similar. However, a closer look (see insets) reveals clear distinct behaviors: $c(x,x^{\prime}=0.9)$ exhibits a more oscillatory structure, driven by the learned weighted periodic kernel, whereas $c(x,x^{\prime}=0.3)$ presents a smoother behavior.
@@ 3 Results and Discussions
All simulations are repeated 16 times independently (except engineering problems in Section 3.4, which are repeated 8 times) to ensure the results are representative. Also, all the GPs are trained via the L-BFGS optimizer with a learning rate of $0.01$. For further implementation and optimization details, please see Section A in the Appendix.
@@ 3.1 Benchmark problems and metrics
We consider seven benchmark problems with varying levels of nonlinearity and dimensionality. These include analytic functions adopted from the literature [noack2024unifying, picheny2013benchmark] and real-world datasets derived from engineering applications [chen2024adaptive, deng2023data, camacho_elevators, delve_pumadyn32].
@@
ROM: we consider our in-house dataset derived from a reduced‑order model for multiscale fracture simulations of cast aluminum alloys [deng2023data]. Each sample contains $6$ real-valued features (four microstructural descriptors and two material properties) and a single output corresponding to the toughness of the microstructure. In each repetition, we train the models with $200$ samples randomly drawn from the full dataset and test their performance on the remaining $50$ samples.
Elevators: this dataset [camacho_elevators] captures the behavior of an aircraft elevator control system. It has $18$ continuous inputs representing the aircraft’s current flight state and recent elevator commands, and a single response showing the desired elevator deflection for the next control step. We discard eight of the input features as they do not affect the response and for each repetition we randomly draw $1{,}750$ training samples ($20\%$ of the full training set) and evaluate the performance of the models on $7{,}847$ test samples.
Pumadyn: this dataset [delve_pumadyn32] was derived from a realistic high-fidelity simulation of the dynamics of a Unimation PUMA 560 robot arm. We use the puma32H variant with moderate output noise, where each data point contains $32$ continuous input features including angular positions, velocities, torques and additional dynamic parameters of the robot arms, and a single continuous target corresponding to the angular acceleration of the arm’s sixth link. For each repetition, we randomly draw $2{,}699$ training samples ($60\%$ of the full training set) and use $3{,}692$ test samples to assess the performance of the models.
@@ 3.2 Sensitivity Studies
We evaluate the performance of the models in terms of NRMSE and NNOIS. The results of these studies are summarized in Figure 3, where each row corresponds to a different study. The analyses are assessed on the Analytic II function and the Volcano dataset. We note that the solid lines and shaded bands in panels (a-d) and (m-p) represent the mean and $95\%$ confidence intervals across repetitions, respectively. The red lines in panels (e-h) and (i-l) indicate the first, second and third quartiles across repetitions, while the blue line indicates the mean.
@@
This observation poses an interesting dilemma: while nonstationary kernels are more beneficial in low-to-mid data regimes, they typically rely on a larger number of parameters. This increased flexibility makes them more prone to overfitting compared to stationary kernels, especially when samples are scarce. This trade-off implies that one must constrain the kernel’s complexity (i.e., number of parameters) according to both the input dimensionality and the available dataset size to prevent overfitting.
@@
Figure 4: Our proposed kernel SEEK with (a) 1 and (b) 6 Gaussian kernels as base kernels on the Volcano dataset: Increasing the number of base kernels reduces error and improves confidence, as reflected in the colorbars (same ranges are used in (a) and (b) to facilitate direct comparison). Black dots represent the training data points.
@@ 3.3 Comparative Studies
In our studies, we noticed that more complex nonstationary kernels (with higher number of parameters) generally require more data to outperform the Gaussian kernel in high-dimensional, sparse settings where effectively learning input-dependent variations becomes especially challenging. We attribute this to the fact that, while these nonstationary kernels introduce additional flexibility to capture such local variations, they also incur a higher risk of overfitting when data are limited. In contrast, the Gaussian kernel imposes a globally smooth prior, acting as an effective regularizer. This observation aligns with prior works [remes2018neural, risser2020bayesian, noack2023exact], which combine flexible nonstationary kernels with sparse approximation techniques to allow the GP to leverage large datasets in high dimensions.
@@
3.4 Engineering Problems
In this section, we evaluate SEEK on three challenging, high-dimensional engineering regression problems (described in Section 3.1): our in‑house ROM dataset, and the popular Elevators and Pumadyn benchmarks. For these three problems, we compare SEEK against the standard Gaussian kernel. The Elevators and Pumadyn datasets have been widely tested in the literature, which allows us to also compare SEEK’s performance against regression models other than GPs [moura2021mine, zhou2023synthetic, sigrist2021ktboost, waxman2024dynamic, lay2012artificial]. These studies aim to show that SEEK provides more accurate mean predictions and improved uncertainty quantification in realistic engineering datasets.
Figure 6: Engineering Problems: We compare the performance of GP using SEEK versus a standard Gaussian kernel on three high-dimensional engineering problems.
Similar to Section 3.3, we parameterize the weight function $\boldsymbol{w}(\boldsymbol{x})$ and bias function $\boldsymbol{b}(\boldsymbol{x})$ with two separate feed-forward NNs. Each network comprises two hidden layers, with (4, 2) neurons for ROM, (8, 4) for Elevators, and (16, 8) for Pumadyn. We chose these layer widths to roughly scale with input dimensionality while keeping the total parameter count reasonable and manageable.
Figure 6 summarizes the results where we observe that in the ROM dataset SEEK achieves roughly a two-times improvement in terms of NRMSE and NNOIS across repetitions. In the Elevators dataset, SEEK also consistently outperforms the Gaussian kernel, achieving a marginally lower NRMSE while reducing the NNOIS by over three times. However, we observe a different trend in the Pumadyn dataset: the single-Gaussian kernel yields an slightly lower NRMSE, but SEEK still delivers an approximately three-fold improvement in terms of NNOIS. These results highlight, once again, SEEK’s capability in providing more reliable and informative uncertainty estimates, which is particularly valuable in engineering design applications where trustworthy UQ is critical.
Finally, we highlight that across all our experiments we have found that GP models incorporating SEEK tend to converge more consistently and in fewer iterations than those using conventional kernels. In particular, on the Pumadyn  problem, we observed that the single-Gaussian kernel failed to converge under the L-BFGS optimizer and only produced the results shown in Figure 6 when trained with AdamW at an unusually high learning rate of $1.0$ to escape local minima. In contrast, GPs with SEEK converged smoothly under the standard optimizer settings employed throughout our studies. This observation holds across a variety of datasets and kernel settings, suggesting that reasonably increased kernel flexibility can not only boost predictive performance but also improve optimization stability and efficiency.